OpenStudy (annebelleschneiders):

solve 3^(2x)=7^(x-1)

zepdrix (zepdrix):

\[\Large\rm 3^{2x}=7^{x-1}\]Hey Annabelle :) See how our x's are stuck up in the exponent? Let's use logarithm to deal with that. There a log rule that will help get them out of that place.

zepdrix (zepdrix):

Let's take the natural log of each side,\[\large\rm \ln\left(3^{2x}\right)=\ln\left(7^{x-1}\right)\]

zepdrix (zepdrix):

Here is the log rule that I have in mind:\[\large\rm \color{orangered}{\log(a^b)=b\cdot \log(a)}\]Do you understand how that will help us?

OpenStudy (annebelleschneiders):

would the first step to multiply each side to balance the logs?

zepdrix (zepdrix):

Hmmm, Im not sure what you mean :o

zepdrix (zepdrix):

The first step was taking natural log of each side. Next we apply our log rule to each side.

OpenStudy (anonymous):

\[\left( 3^2 \right)^x=7^x*7^{-1},9^x=\frac{ 7^x }{ 7 },\frac{ 9^x }{ 7^x }=\frac{ 1 }{ 7 }\] \[\left( \frac{ 9 }{ 7 } \right)^x=\frac{ 1 }{ 7 }\] x[log9-log7]=log1-log7 x=?